# NeuronSeek: On Stability and Expressivity of Task-driven Neurons

### v1

@@ IV Super-expressive Property Theory of Task-driven Neuron
Remark 2. Our construction saves the parametric complexity. In [18], though the number of parameters remains unchanged, the values of parameters $\theta^{*}$ increase as long as $n$ increases. Therefore, the entire parametric complexity goes up. In contrast, in our construction, the parameter value of $T$ is also fixed for different $(k,f(\tilde{x}_{k}))$. Yet, $T$ is copied and composed $m(k)$ times. The overall parametric complexity remains unchanged. What increases herein is computational complexity. Therefore, we call for brevity our construction “super-super-expressive”.
TABLE II: Regression results (MSE) on synthetic data. $\delta$ represents the noise level.
Ground Truth Formula $\delta=0\%$ $\delta=10\%$ $\delta=20\%$ Time (s)
SR TD SR TD SR TD SR TD
$2\mathbf{x}$ 0.0013 0.0008 0.0108 0.0120 0.0232 0.0138 13.94 4.46
$1.5(x\odot x)$ 0.0007 0.0006 0.0163 0.0095 0.0191 0.0166 12.75 6.69
$(\mathbf{x}\odot^{2}\mathbf{x})-1.5(\mathbf{x}\odot x)+2$ 0.0235 0.0028 0.2489 0.0102 0.1905 0.1706 28.77 7.73
$(\mathbf{x}\odot^{2}\mathbf{x})-1.5(\mathbf{x}\odot\mathbf{x})+3\mathbf{x}$ 0.0016 0.0012 0.0139 0.0101 0.0147 0.0116 22.01 8.82
$(\mathbf{x}\odot^{2}\mathbf{x})+(\mathbf{x}\odot\mathbf{x})$ 0.0095 0.0067 0.0116 0.0086 0.0495 0.0064 15.16 5.80
$4(\mathbf{x}\odot^{3}\mathbf{x})$ 0.0011 0.0002 0.0230 0.0165 0.0561 0.0193 24.77 9.44
$2(\mathbf{x}\odot^{3}\mathbf{x})+(\mathbf{x}\odot^{2}\mathbf{x})$ 0.0055 0.0014 0.0125 0.0104 0.0416 0.0109 15.80 8.21
$(\mathbf{x}\odot^{4}\mathbf{x})$ 0.0102 0.0062 0.0134 0.0111 0.0170 0.0126 13.30 8.24
@@
V Comparative Experiments on NeuronSeek
This section presents a comprehensive experimental evaluation of our proposed method and the previous NeuronSeek-SR [12]. First, we systematically demonstrate the superior performance and enhanced stability of NeuronSeek-TD compared to the previous NeuronSeek-SR. The results demonstrate that pursuing stability in NeuronSeek-TD does not hurt the representation power of the resultant task-driven neurons and even present better results.
V-A Experiment on Synthetic Data
Here, we adopt the synthetic data to compare the performance of NeuronSeek-TD and NeuronSeek-SR. The synthetic data are generated by adding noise perturbations to the known ground-truth functions. We then measure the regression error (MSE) of the formulas of NeuronSeek-SR and NeuronSeek-TD on the synthetic data.
V-A1 Experimental settings
Eight multivariate polynomials are constructed as shown in Table II. For simplicity and conciseness, we prescribe:
$\sum_{i=1}^{d}rx_{i}^{s}=r\cdot\underbrace{(\mathbf{x}\odot\cdots\odot\mathbf{x})^{\top}}_{s\text{ times}}=r\cdot(\mathbf{x}\odot^{s}\mathbf{x})^{\top},$
where $\bm{r}=[r,\ldots,r]\in\mathbb{R}^{d}$ denotes a constant vector, $\bm{x}=[x_{1},\ldots,x_{d}]$ represents the input vector, and $\odot$ denotes the Hadamard product. The elements of $\bm{x}$ are uniformly sampled from the interval $[-1,1]$. We set $d=30$ to simulate multi-dimensional inputs.
To evaluate the noise robustness of the proposed methods, we introduce three noise levels through
$\bar{y}=y+\delta\cdot\mathrm{RMS}(y)\cdot\bm{\eta},$
where $y$ denotes the noise-free polynomial output, $\bar{y}$ is the noisy observation, $\bm{\eta}\sim\mathcal{N}(\bm{0},\bm{I})$ is a standard normal random vector, $\delta$ controls the relative noise level, and $\mathrm{RMS}(y)$ represents the root mean square (RMS) of the data:
$\mathrm{RMS}(y)=\sqrt{\frac{1}{N}\sum_{i=1}^{N}y_{i}^{2}}.$
Here, $N=1,000$ specifies the number of data points. For NeuronSeek-SR, we configure a population size of 1,000 with the maximum of 50 generations. NeuronSeek-TD is optimized using Adam for 50 epochs. The noise levels are set to $\delta\in\{0\%,10\%,20\%\}$ respectively. We quantify the regression performance using the mean squared error (MSE) of the discovered neuronal formulae. Lower MSE values indicate better preservation of the underlying data characteristics.
Figure 8: The uniqueness experiment on phoneme data.
V-A2 Experimental results
The performance comparison on the synthetic data is presented in Table II. Our proposed NeuronSeek-TD demonstrates consistent superiority over NeuronSeek-SR across all generated data, with particularly notable improvements on the data derived from the function $(\mathbf{x}\odot^{2}\mathbf{x})-1.5(\mathbf{x}\odot\mathbf{x})+2$. This performance advantage remains robust under varying noise levels.
To assess computational efficiency, we measure the average execution time for both methods. NeuronSeek-TD achieves substantial computational savings, requiring less than half the execution time of the SR-based approach while maintaining comparable solution quality. These results collectively demonstrate that our method achieves both superior accuracy and greater computational efficiency.
V-B Experiments on Uniqueness and Stability
Here, we demonstrate that NeuronSeek-SR exhibits instability in terms of resulting in divergent formulas under minor perturbations in initialization, while NeuronSeek-TD maintains good consistency.
V-B1 Experimental settings
Our experimental protocol consists of two benchmark datasets: the phoneme dataset (a classification task classifying nasal and oral sounds) and the airfoil self-noise dataset (a regression task modeling self-generated noise of airfoil). These public datasets are chosen from the OpenML website to represent both classification and regression tasks, ensuring a comprehensive evaluation across different problem types. For each dataset, we introduce four levels of Gaussian noise ($\sigma\in\{0.01,0.025,0.05,0.1\}$) to simulate real-world perturbations. We train 10 independent instances of each method under identical initialization. For comparison, NeuronSeek-SR based on GP uses a population of 3,000 candidates evolved over 20 generations, whereas NeuronSeek-TD methods—employing CP decomposition and Tucker decomposition—rely on a model optimized via Adam over 20 epochs.
V-B2 Experimental results
@@
The results of phoneme and airfoil self-noise datasets are depicted in Figures 7 and 8, respectively. From the epoch-wise diversity (Figure 7 (a) and 8 (a)), we have several important observations: Firstly, GP-based SR demonstrates a rapid increase in the number of discovered formulas within each epoch. After training, it produces more unique formulas compared to TD-based methods (CP and Tucker). This indicates that SR is sensitive to initialization conditions. Secondly, as noise levels increase, the GP method discovers nearly 10 distinct formulas across 10 initializations. With higher noise, every initialization produces a completely different formula structure. These behaviors highlight the instability of GP-based SR when dealing with noisy data, as minor perturbations in initialization or data noise can lead to substantially different symbolic expressions.
TABLE III: Comparison of the network built by neurons using formulas generated by NeuronSeek-SR and NeuronSeek-TD. The number $h_{1}-h_{2}-\cdots-h_{k}$ means that this network has $k$ hidden layers, each with $h_{k}$ neurons. For example, 5-3-1 means this network has two hidden layers with 5 and 3 neurons respectively in each layer.
Datasets Instances Features Classes Metrics Structure Test results
TN NeuronSeek-TD NeuronSeek-SR NeuronSeek-TD
California housing 20640 8 continuous MSE 6-4-1 6-3-1 0.0720 $\pm$ 0.0024 0.0701 $\pm$ 0.0014
house sales 21613 15 continuous 6-4-1 6-4-1 0.0079 $\pm$ 0.0008 0.0075 $\pm$ 0.0004
airfoil self-noise 1503 5 continuous 4-1 3-1 0.0438 $\pm$ 0.0065 0.0397 $\pm$ 0.0072
diamonds 53940 9 continuous 4-1 4-1 0.0111 $\pm$ 0.0039 0.0102 $\pm$ 0.0015
abalone 4177 8 continuous 5-1 4-1 0.0239 $\pm$ 0.0024 0.0237 $\pm$ 0.0027
Bike Sharing Demand 17379 12 continuous 5-1 5-1 0.0176 $\pm$ 0.0026 0.0125 $\pm$ 0.0146
space ga 3107 6 continuous 2-1 2-1 0.0057 $\pm$ 0.0029 0.0056 $\pm$ 0.0030
Airlines DepDelay 8000 5 continuous 4-1 4-1 0.1645 $\pm$ 0.0055 0.1637 $\pm$ 0.0052
credit 16714 10 2 ACC 6-2 6-2 0.7441 $\pm$ 0.0092 0.7476 $\pm$ 0.0092
heloc 10000 22 2 18-2 17-2 0.7077 $\pm$ 0.0145 0.7089 $\pm$ 0.0113
electricity 38474 8 2 5-2 5-2 0.7862 $\pm$ 0.0075 0.7967 $\pm$ 0.0087
phoneme 3172 5 2 8-5 7-5 0.8242 $\pm$ 0.0207 0.8270 $\pm$ 0.0252
MagicTelescope 13376 10 2 6-2 6-2 0.8449 $\pm$ 0.0092 0.8454 $\pm$ 0.0079
vehicle 846 18 4 10-2 10-2 0.8176 $\pm$ 0.0362 0.8247 $\pm$ 0.0354
Oranges-vs-Grapefruit 10000 5 2 8-2 7-2 0.9305 $\pm$ 0.0160 0.9361 $\pm$ 0.0240
eye movements 7608 20 2 10-2 10-2 0.5849 $\pm$ 0.0145 0.5908 $\pm$ 0.0164
@@
Furthermore, regarding cumulative diversity (Figures 7(b) and 8(b)), we can observe that GP-based SR constantly discovers an increasing number of different formulas, while TD methods tend to stabilize during training. This indicates that the GP algorithm struggles to converge to a consistent symbolic formula. In contrast, the asymptotic behavior exhibited by TD-based methods demonstrates their superior stability and convergence properties when identifying symbolic expressions. Moreover, the CP-based method suggests the best resilience against initialization variability.
V-C Superiority of NeuronSeek-TD over NeuronSeek-SR
@@
To validate the superior performance of NeuronSeek-TD over NeuronSeek-SR, we conduct experiments across 16 benchmark datasets—8 for regression and 8 for classification—selected from scikit-learn and OpenML repositories. First, we use NeuronSeek-TD and NeuronSeek-SR simultaneously to search for the optimal formula. Second, we construct fully connected networks (FCNs) with neurons taking these discovered formulas as their aggregation functions. To verify the ability of NeuronSeek-TD, the structures of neural networks are carefully set, such that NeuronSeek-TD has equal or fewer parameters than NeuronSeek-SR.
@@
Comprehensive implementation details and test results are provided in Table III. For regression tasks, the dataset is partitioned into training and test sets in an 8:2 ratio. The activation function is ReLU with MSE as the loss function and RMSProp as the optimizer. For classification tasks, the same train-test split is applied, using the Sigmoid activation function. The loss function and optimizer are set to CE and Adam, respectively. All experiments are repeated 10 times and reported by mean and standard deviation, with MSE for regression and accuracy for classification.
The regression results in Table III demonstrate that NeuronSeek-SR exhibits larger fitting errors compared to NeuronSeek-TD, along with higher standard deviations in MSE values. A notable example is the airfoil self-noise dataset, where NeuronSeek-TD outperforms NeuronSeek-SR by a significant margin, achieving a 9% reduction in MSE. Furthermore, NeuronSeek-TD shows superior performance on both the California housing and abalone datasets while utilizing fewer parameters. These results collectively indicate that NeuronSeek-TD possesses stronger regression capabilities than NeuronSeek-SR.
This performance advantage extends to classification tasks, where NeuronSeek-TD consistently achieves higher accuracy than NeuronSeek-SR. Particularly in the HELOC and orange-vs-grapefruit datasets, NeuronSeek-TD attains better classification performance with greater parameter efficiency. The observed improvements across both regression and classification tasks suggest that NeuronSeek-TD offers superior modeling capabilities compared to NeuronSeek-SR.
V-D Comparison of Tensor Decomposition Methods

### v2

@@ 4 Super-expressive Property Theory of Task-driven Neurons
Remark 2. Our construction reduces the parametric complexity. In [54], though the number of parameters remains unchanged, the values of parameters $\theta^{*}$ increase as long as $n$ increases. Therefore, the overall parametric complexity still goes up, as it consumes more digits. In contrast, in our construction, the parameter value of $T$ is also fixed for different $(k,f(\tilde{x}_{k}))$. Yet, $T$ is copied and composed $m(k)$ times. The overall parametric complexity remains unchanged. What increases herein is computational complexity. Therefore, we call for brevity our construction “super-super-expressive”.
5 Analysis Experiments of NeuronSeek
5.1 Experiment on Synthetic Data
In this section, we compare the performance of our proposed NS-TD with former NS-SR[17] and traditional symbolic regression methods on synthetic data. We select representative benchmark methods including classic symbolic regression [52], EQL [45], and the state-of-the-art method MetaSymNet [37]. Synthetic dataset is generated by the functions in Table 2. For each task, input vectors $\mathbf{x}\in\mathbb{R}^{d}$ are sampled from a standard normal distribution $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{d})$, and the target $\mathbf{y}$ is standardized to zero mean and unit variance. Crucially, regarding the function coefficients, while most are sampled from a normal distribution to ensure diversity, we deliberately assign high-contrast values (e.g., 10 vs. 0.5) in specific cases. This design simulates signal masking scenarios, thereby increasing the difficulty of capturing weak structural information amidst dominant features.
Table 2: Synthetic Dataset Generation Formulas. Comparisons of Pure, Interaction, and Hybrid modes. Coefficients $c_{i}$ are sampled from $\mathcal{N}(0,1)$.
ID Pure Mode Interact Mode Hybrid Mode
$f(\mathbf{x})\sim P_{k}$ $f(\mathbf{x})\sim I_{k}$ $f(\mathbf{x})\sim P_{k}+I_{k}$
0 $c\cdot P_{2}(\mathbf{x})$ $c\cdot I_{2}(\mathbf{x})$ $c_{1}P_{2}(\mathbf{x})+c_{2}I_{2}(\mathbf{x})$
1 $c\cdot P_{3}(\mathbf{x})$ $c_{1}I_{2}(\mathbf{x})+c_{2}I_{3}(\mathbf{x})$ $c_{1}P_{1}(\mathbf{x})+c_{2}I_{2}(\mathbf{x})$
2 $10P_{1}(\mathbf{x})+0.5P_{2}(\mathbf{x})$ $8I_{2}(\mathbf{x})+0.5I_{3}(\mathbf{x})$ $5P_{3}(\mathbf{x})+0.5I_{2}(\mathbf{x})$
3 $5P_{2}(\mathbf{x})+0.5P_{4}(\mathbf{x})$ $c_{1}I_{2}(\mathbf{x})+c_{2}I_{4}(\mathbf{x})$ $0.5P_{2}(\mathbf{x})+5I_{3}(\mathbf{x})$
4 $c\cdot P_{5}(\mathbf{x})$ $c\cdot I_{4}(\mathbf{x})$ $c_{1}P_{2}(\mathbf{x})+c_{2}I_{2}(\mathbf{x})+c_{3}I_{3}(\mathbf{x})$
Note: $P_{k}(\mathbf{x})=\sum_{j=1}^{d}x_{j}^{k}$ denotes the element-wise pure polynomial term. $I_{k}(\mathbf{x})=\prod_{m=1}^{k}(\mathbf{w}_{m}^{\top}\mathbf{x})$ denotes the interaction term of order $k$, where $\mathbf{w}_{m}\sim\mathcal{N}(0,1/\sqrt{d})$.
The experimental pipeline is structured into two distinct stages: structure discovery and structure evaluation. In the first stage, all methods operate under a strict time budget of 60 seconds to search for the underlying regression formula of a given dataset. Subsequently, we extract the symbolic structure of the discovered formula and proceed to the second stage, where we re-optimize the constants from scratch to evaluate the model’s representational capacity. This two stage method is designed to ensure that we solely evaluate the capacity of searched structure. For our proposed NS-TD, we set the tensor rank to 8, which provides sufficient capacity for the synthetic functions, relying on the model’s sparsity mechanism to prune redundant components. The regularization coefficient is set to 0.05, a value empirically selected to balance reconstruction fidelity with structural sparsity. For genetic programming-based methods (Standard-SR and TN-SR), following the configuration in [17], we set the population size to 1,000, with crossover, mutation, and reproduction probabilities of 30%, 60%, and 10%, respectively; an elitism strategy is applied to select the top 5% of individuals for the next generation. For MetaSymNet [37], we adopt an entropy loss regularization of 0.2 and a learning rate of 0.01. The hyperparameter settings for EQL follow those in [45]. For data preparation, we generate 2,500 samples per function, split into training and test sets with an 80/20 ratio.
We report the average MSE across five distinct functions in Table 3. NS-TD demonstrates a decisive advantage over competing symbolic regression baselines. Across all functional modes and dimensionalities, our method consistently achieves the lowest test MSE, validating its superior capability in identifying mathematical structures that accurately approximate high-dimensional data manifolds. In contrast, while the tensor network-based NS-SR outperforms traditional discrete search methods like PySR and MetaSymNet in the Pure Mode, it exhibits a significant performance deterioration in the Interact and Hybrid modes. This performance gap stems from NS-SR’s architectural reliance on univariate transformations, which precludes the modeling of cross-feature interactions, thereby failing to capture the coupled dynamics inherent in the latter two scenarios. Moreover, NS-TD displays remarkable stability as the input dimension scales from d=10 to d=100. Unlike baseline methods which suffer from convergence difficulties, our approach maintains consistently low error rates, confirming its scalability and robustness in handling structure discovery tasks with high-dimensional data, which is common in deep learning scenarios. Moreover, the Standard Error of the Mean (SEM) across different trials are reported in Table 911 in the Appendix.
Table 3: Performance Analysis on Synthetic Data. We evaluate the reconstruction accuracy (Mean MSE) across Pure, Interact, and Hybrid modes under varying input dimensions ($d\in\{10,30,50,100\}$). Best results are highlighted in bold.
Method Pure Mode Interact Mode Hybrid Mode
($\sum x^{k}$) (Vector Interact) (Mixture)
10 30 50 100 10 30 50 100 10 30 50 100
SR (PySR) 1.1612 1.1493 1.2325 1.2481 0.8754 1.0042 0.9776 0.9763 1.0635 0.9821 1.0764 1.1018
EQL 0.4524 0.5342 0.5705 0.6923 0.3875 0.4862 0.5614 0.4521 0.3126 0.4113 0.4237 0.4472
NS-SR (VSR) 0.2324 0.2541 0.2825 0.2953 0.6972 0.7734 0.7326 0.8145 0.7783 0.7824 0.8281 0.8685
MetaSymNet 0.8324 0.8541 0.9825 0.9843 0.8632 1.0804 1.0193 0.9942 1.1345 0.9526 1.1093 1.1334
NS-TD 0.0543 0.0632 0.0854 0.0891 0.0512 0.0583 0.0724 0.0945 0.0423 0.0621 0.0735 0.1112
Convergence and Efficiency. To further investigate the optimization dynamics, we analyze the convergence trajectory of all methods on the most challenging task (Hybrid Mode, $d=100$), as illustrated in Fig. 7. We can observe that our method NS-TD (Red line) converges to the optimal solution faster than evolutionary baselines. This confirms that our differentiable architecture search effectively bypasses the combinatorial search space that hinders methods like NS-SR and SR. Though EQL stabilizes at a plateau at 5.5 seconds, quicker than NS-TD (approximately at 10s), this higher plateau indicates that EQL is stuck in local optima.
Figure 7: Convergence Efficiency Analysis (Hybrid Mode, $d=100$).
Why Traditional Symbolic Regression Fails? The failure of baseline methods stems from two fundamental limitations. First, the combinatorial explosion in high-dimensional search spaces prevents discrete search algorithms from converging into a good solution within practical time budgets. Second, and more critically, traditional SR tends to generate structurally irregular formulas that are mathematically incompatible with neural architectures that favor homogeneous formulas.
Due to the prohibitive length and structural heterogeneity of baseline outputs, we present a representative case study in Table 4. As illustrated, baseline methods suffer from consistently observed structural pathologies across high-dimensional tasks: SR generates index-dependent heterogeneous terms; EQL results in excessive density; and MetaSymNet collapses into local features. These heterogeneous structures violate the homogeneity required for tensor operations, meaning they cannot be efficiently vectorized or broadcasted across features. Consequently, unlike the compact forms discovered by NS-TD, these irregular baselines are computationally intractable to serve as scalable aggregation functions for dense deep learning layers.
Deep learning tasks predominantly involve high-dimensional data representations. Through our experiments simulating these high-dimensional generation processes, NS-TD has demonstrated exceptional efficiency in identifying formulas that accurately capture the underlying data patterns. Crucially, in contrast to traditional symbolic regression approaches that yield irregular and computationally intractable expressions, the formulas discovered by NS-TD exhibit the structural homogeneity and compactness required for effective neuronal aggregation. This confirms that NS-TD uniquely satisfies the dual objectives of high reconstruction fidelity and structural suitability for constructing task-based neurons.
Table 4: Case Study: Discovered Structures on Hybrid Mode (Formula ID:0, $d=100$). We illustrate the symbolic forms identified by different methods. While NS-TD successfully recovers the homogeneous vectorized structure, baselines generate heterogeneous or dense formulas that are ill-suited for neuronal aggregation.
Method Discovered Structure
Ground Truth $P_{2}(\mathbf{x})+I_{2}(\mathbf{x})$
NS-TD (Ours) $P_{2}(\mathbf{x})+I_{2}(\mathbf{x})$
NS-SR $P_{1}(\mathbf{x})+P_{2}(\mathbf{x})$
SR $\frac{x_{12}+x_{19}}{x_{16}}+x_{25}^{2}-0.05x_{12}\dots$
EQL Mixture of 79 active terms ($\sum w_{i}\phi_{i}(\mathbf{x})$)
MetaSymNet $\log(\exp(x_{82}))+x_{45}^{2}$
5.2 Experiments on Uniqueness and Stability
Here, we demonstrate that NS-SR exhibits instability in terms of resulting in divergent formulas under minor perturbations in initialization, while NS-TD maintains good consistency.
5.2.1 Experimental Settings
Our experimental protocol consists of two benchmark datasets: the phoneme dataset (a classification task classifying nasal and oral sounds) and the airfoil self-noise dataset (a regression task modeling self-generated noise of airfoil). These public datasets are chosen from the OpenML website to represent both classification and regression tasks, ensuring a comprehensive evaluation across different problem types. For each dataset, we introduce four levels of Gaussian noise ($\sigma\in\{0.01,0.025,0.05,0.1\}$) to simulate real-world perturbations. We train 10 independent instances of each method under identical initialization. For comparison, NS-SR based on GP uses a population of 3,000 candidates evolved over 20 generations, whereas NS-TD methods—employing CP decomposition and Tucker decomposition—rely on a model optimized via Adam over 20 epochs.
  • 1.
  • 2.
5.2.2 Experimental Results
The results of phoneme and airfoil self-noise datasets are depicted in Figs. 8 and 9, respectively.
(a) Epoch-wise diversity
(b) Cumulative diversity (average per initialization)
@@ 5.2.2 Experimental Results
(a) Epoch-wise diversity
(b) Cumulative diversity (average per initialization)
Figure 9: Uniqueness experiment on phoneme data.
@@
From the epoch-wise diversity (Fig. 8 (a) and Fig. 9 (a)), we have several highlights: Firstly, GP-based SR demonstrates a rapid increase in the number of discovered formulas within each epoch. After training, it produces more unique formulas compared to TD-based methods (CP and Tucker). This indicates that SR is sensitive to initialization conditions. Secondly, as noise levels increase, the GP method discovers nearly 10 distinct formulas across 10 initializations. With higher noise, every initialization produces a completely different formula structure. These behaviors highlight the instability of GP-based SR when dealing with noisy data, as minor perturbations in initialization or data noise can lead to substantially different symbolic expressions.
@@
Furthermore, regarding cumulative diversity (Fig. 8(b) and Fig. 9(b)), we can observe that GP-based SR constantly discovers an increasing number of different formulas, while TD methods tend to stabilize during training. This indicates that the GP algorithm struggles to converge to a consistent symbolic formula. In contrast, the asymptotic behavior exhibited by TD-based methods demonstrates their superior stability and convergence properties when identifying symbolic expressions. Moreover, the CP-based method suggests the best resilience against initialization variability.
5.3 Superiority of NS-TD
@@
To validate the superior performance of NS-TD, we conduct experiments across 16 benchmark datasets—8 for regression and 8 for classification—selected from scikit-learn and OpenML repositories. This experiment is also conducted in 2 stages. First, we use NS-TD and NS-SR simultaneously to search for the optimal formula. Second, we construct fully connected networks (FCNs) with neurons taking these discovered formulas as their aggregation functions. We evaluate the MSE on regression tasks and accuracy on classification tasks. For comparison, we select Kolmogorov-Arnold Network(KAN) [38] and multi-layer perceptron(MLP) as baselines.
@@
For regression tasks, the dataset is partitioned into training and test sets in an 8:2 ratio. The activation function is ReLU with MSE as the loss function and RMSProp as the optimizer. For classification tasks, the same train-test split is applied, using the Sigmoid activation function. The loss function and optimizer are set to CE and Adam, respectively. All experiments are repeated 10 times and reported by mean. We also report #parameters and #FLOPs to show the computation cost of different methods.
[…]
Table 5: Comprehensive Comparison on Benchmark Datasets. We evaluate methods on both regression (MSE $\downarrow$) and classification (Accuracy $\uparrow$) tasks. To assess efficiency, we also report FLOPs and Params. Note that NS-TD achieves comparable or better performance with significantly lower computational costs.
Dataset MLP KAN NS-SR NS-TD (Ours)
Perf. FLOPs Param Perf. FLOPs Param Perf. FLOPs Param Perf. FLOPs Param
Regression Tasks (Metric: MSE $\downarrow$)
California Housing 0.0890 1568 833 0.1017 19544 7056 0.1135 3104 1601 0.1005 1568 1857
House Sales 0.0077 752 401 0.0073 10551 3384 0.0075 1488 769 0.0075 2224 1137
Airfoil 0.0326 152 89 0.0220 3017 684 0.0712 152 89 0.0209 152 393
Diamonds 0.0053 560 305 0.0040 8241 2520 0.0043 1104 577 0.0045 560 897
Abalone 0.0249 200 113 0.0232 3740 900 0.0243 392 209 0.0236 584 305
Bike Sharing 0.0250 656 353 0.0086 6460 1872 0.0081 1296 673 0.0065 1936 993
Space GA 0.0064 168 97 0.0056 3258 756 0.0063 328 177 0.0056 488 257
Airlines Delay 0.1605 848 465 0.1563 11997 3816 0.1602 1680 881 0.1560 2512 1297
Classification Tasks (Metric: Accuracy $\uparrow$)
Credit 0.7341 608 330 0.7431 3474 864 0.7441 1184 618 0.7471 1760 906
Heloc 0.7107 432 230 0.7027 7186 1944 0.7117 848 438 0.7077 432 806
Electricity 0.7762 544 298 0.7962 2992 720 0.7932 1056 554 0.7862 1568 810
Phoneme 0.7742 448 250 0.8292 6845 2016 0.7942 864 458 0.8242 448 778
MagicTelescope 0.8444 608 330 0.8479 8770 2736 0.8389 1184 618 0.8495 1184 1226
Vehicle 0.8106 896 476 0.7576 12138 4032 0.7676 1728 892 0.8176 2560 1308
Oranges-vs-Grape 0.9105 448 250 0.9302 6845 2016 0.9235 864 458 0.9344 448 250
Eye Movements 0.5699 400 214 0.5819 6704 1800 0.5899 784 406 0.5940 400 758
5.4 Comparison of Tensor Decomposition Methods